OpenStudy (surryyy):

Select the equations that are equivalent to the equation y-4=5(x+2) . Select all that apply. y=2x+5 y=5x+14 5x-y=-14 2x+y=14

imqwerty (imqwerty):

y-4=5(x+2) y-4=5x+10 y=5x+14 can u look at options and tell which equations look similar to this one :)

OpenStudy (surryyy):

lol b?

imqwerty (imqwerty):

yes :) and what abt c?

OpenStudy (surryyy):

so my answers would be b and c?

imqwerty (imqwerty):

5x-y=-14 add 14 to both sides u get- 5x-y+14=-14+14 5x+14-y=0 add y to both sides u get- 5x-y+14+y=y y=5x+14 so c is also similar and what abt a

OpenStudy (surryyy):

a isnt similar

imqwerty (imqwerty):

yes :) hehe :D so ur answers are b and c :)

OpenStudy (surryyy):

lol thnx

imqwerty (imqwerty):

welcome :) ur ss went up by 1

OpenStudy (surryyy):

haha cx

imqwerty (imqwerty):

cx